OpenStudy (anonymous):

An equation is shown below: 4(2x – 5) = 4 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation? Show your work. (6 points)

OpenStudy (anonymous):

@Kash_TheSmartGuy @Pawanyadav

OpenStudy (johnweldon1993):

Any idea here? :)

OpenStudy (johnweldon1993):

\[\large 4(2x - 5) = 4\] What would be the first step here?

OpenStudy (pawanyadav):

2x-5=4/4 2x-5=1 2x=6 x=3 Only one solution as equation of degree One.